OpenStudy (goldps):

3x^3-5x^2-4x-2 divided by 2x^2+x

OpenStudy (goldps):

i'm stuck on the first bit, how do i get 3x^3 to cancel with 2x^2?

OpenStudy (anonymous):

\[\frac{3x^3}{2x^2}=\frac{3}{2}x\]